The 2nd and 5th terms of a Geometric Progression are 7 and 56 respectively . (a) find the common ratio . (b) the first term . (c) sum of the first 5 terms
Answer by Boreal(15235)

An=A1*r^(n-1)
7=A1r
56=A1r^4
56=8A1r
therefore
8A1r=A1r^4
divide by A1 then divide by r
8=r^3
r=2 common ratio
first term is 7/2 or 3.5
sum of the first five terms
S=A1(1-r^n)/(1-r)
=3.5(1-2^5)(-1)
=3.5*31=108.5
terms are 3.5,7,14,28,56=108.5